OpenStudy (vannieduval):

Jamie is painting her house. Each room takes \dfrac{1}{2} ​2 ​ ​1 ​​ start fraction, 1, divided by, 2, end fraction of a liter of paint. If Jamie has 333 liters of paint, how many rooms can she paint?

OpenStudy (mathstudent55):

I don't understand the problem. Can you post a picture of the problem or rewrite it?

OpenStudy (vannieduval):

@vanilla_ice03 @Aaron96

OpenStudy (aaron96):

how many rooms does the house have

OpenStudy (vannieduval):

it does not say i copyed and pasted the quetion

OpenStudy (mathstudent55):

Each room takes \(\dfrac{1}{2} \) liter of paint?

OpenStudy (mathstudent55):

Divide the total amount of paint by the amount needed for 1 room.

OpenStudy (aaron96):

To be honest, I really don't know the answer to this question. If I knew the number of rooms that the house had then I could probably tell you.

OpenStudy (vannieduval):

0.4

OpenStudy (mathstudent55):

The problem does not give the number of rooms because the question is how many rooms can you paint with the given amount of paint. The problem here is not being able to read the post. The question is simple. If you know how much paint you need for 1 room, and you know how much total paint you have, divide the total amount of paint by the paint you need for 1 room to find out how many rooms you have paint for.

OpenStudy (vannieduval):

is it 6

OpenStudy (vannieduval):

On Monday, Daniel and Emily decided to see how fast they could sprint 400400400 meters. They asked their friend Ben to time them with a stopwatch. After 4.534.534, point, 53 minutes, Ben agreed to time the runners. Daniel sprinted first and ran 400400400 meters in 64.1864.1864, point, 18 seconds. When it was Emily's turn, she sped off and completed the run in 61.0661.0661, point, 06 seconds. How much faster was Emily than Daniel in seconds?

OpenStudy (vannieduval):

@Mythorius @lia22 @Aaron96

OpenStudy (mathstudent55):

Have you read your questions? In the first question, is there 3 liters of paint, and the room takes 1/3 liter of paint? In the original post, it says there are 333 liters of paint. If there are 3 liters of paint, then divide 3 by 1/2 to find the answer. \(3 \div \dfrac{1}{2} = \dfrac{3}{1} \times \dfrac{2}{1} = \dfrac{6}{1} = 6\) The paint is enough for 6 rooms.

OpenStudy (mathstudent55):

Read your second question. The numbers are unintelligible. Rewrite the question with numbers that can be understood.